# Neural variational inference for cutting feedback during uncertainty propagation

### v1

@@ C.3 Rational Quadratic Neural Spline Flows
$m\ \leq\ C_{r}\bigl(4K_{\star}+2\bigr)\left(\frac{\bigl(2e^{2H(M)}L_{b}(M)+L_{\mathrm{edge}}(M)\bigr)\,\mathrm{diam}(K_{\eta})}{\varepsilon_{\mathrm{par}}}\right)^{d_{\eta}}$
$\times\log\!\left(\frac{\bigl(2e^{2H(M)}L_{b}(M)+L_{\mathrm{edge}}(M)\bigr)\,\mathrm{diam}(K_{\eta})}{\varepsilon_{\mathrm{par}}}\right),$
@@
$K_{\star}=\Bigl\lceil\frac{L_{b}(M)M}{\varepsilon_{d}}\Bigr\rceil,\qquad\varepsilon_{d}=\frac{\varepsilon}{4\,M\,e^{H(M)}},\qquad\varepsilon_{\mathrm{par}}=\min\!\Big\{\frac{\varepsilon}{2},\ \frac{\varepsilon}{4\,M\,e^{H(M)}}\Big\}.$
Appendix D Full Bayes and exact cut in misspecified prior
@@
Full Bayes posterior
@@
Appendix E Estimating Cause-Specific Mortality Fractions in Mozambique for Child population
@@
Figure 6 presents the estimated densities of the cause-specific mortality fractions (CSMFs) for the child population, consistent with those obtained for the neonate population. The sequential Bayes posteriors differ markedly from the corresponding cut-posteriors. Among the cut-Bayes approaches, NeVI-Cut exhibits close alignment with nested MCMC, whereas the parametric cut, based on a Dirichlet variational family, shows limited flexibility and induces bias in estimation.
@@
Figure 6: Posterior distributions of calibrated cause-specific mortality fractions (CSMFs) for children aged 1-59 months, based on VA-only data collected by the Countrywide Mortality Surveillance for Action (COMSA) program in Mozambique. NeVI-Cut effectively captures complex distributional characteristics, such as multimodality and skewness, achieving performance comparable to nested MCMC.

### v2

@@ Acknowledgements
$m\ \leq\ C_{r}\bigl(4K_{\star}+2\bigr)\left(\frac{\bigl(2e^{2H(M)}\mathsf{L}_{b}(M)+L_{\mathrm{edge}}(M)\bigr)\,\mathrm{diam}(K_{\eta})}{\varepsilon_{\mathrm{par}}}\right)^{d_{\eta}}$
$\times\log\!\left(\frac{\bigl(2e^{2H(M)}\mathsf{L}_{b}(M)+L_{\mathrm{edge}}(M)\bigr)\,\mathrm{diam}(K_{\eta})}{\varepsilon_{\mathrm{par}}}\right),$
@@
$K_{\star}=\Bigl\lceil\frac{\mathsf{L}_{b}(M)M}{\varepsilon_{d}}\Bigr\rceil,\qquad\varepsilon_{d}=\frac{\varepsilon}{4\,M\,e^{H(M)}},\qquad\varepsilon_{\mathrm{par}}=\frac{\varepsilon}{4\,M\,e^{H(M)}}.$
Once again to apply Theorem 1, one has to replace $\epsilon$ by $\varepsilon/O(\tilde{L}_{S}(\tilde{G}(M(\varepsilon)))$ which yields the expressions in the statement of Theorem 3.
Appendix S6 Details of Simulation Experiments
S6.1 Misspecified Prior
We revisit a commonly used illustrative example (Liu et al.,, 2009; Jacob et al.,, 2017; Yu et al.,, 2023) that demonstrates the challenges of full Bayesian inference when there is model misspecification or biased data. This example constitutes a case of a misspecified prior in the downstream analysis. The upstream dataset consists of a small sample $z=(z_{1},\ldots,z_{n_{1}})^{\top}$, where $z_{i}\overset{\text{iid}}{\sim}N(\varphi,1)$. The downstream dataset is a large sample $w=(w_{1},\ldots,w_{n_{2}})^{\top}$, with $w_{i}\overset{\text{iid}}{\sim}N(\varphi+\eta,1)$. This dataset has a mean $\varphi+\eta$, where $\eta$ represents a downstream bias parameter. We use the same priors as in prior research, i.e., $\varphi\sim N(0,\delta_{1}^{-1})$ and $\eta\sim N(0,\delta_{2}^{-1})$, with $\delta_{2}$ set to a large value to reflect high prior confidence in the absence of bias. When the true bias is large, this prior leads to a biased posterior under the full Bayesian model due to the overconfident assumption that $\eta\approx 0$.
This simulation setting illustrates how posterior samples from the small upstream dataset can be leveraged to inform inference on the bias parameter $\eta$ for the large dataset. Following previous studies, we set $n_{1}=100$ and $n_{2}=1000$, with true values $\varphi=0$ and $\eta=1$. The inverse variance parameters are set to $\delta_{1}=1$ and $\delta_{2}=100$. For this setting, the full Bayes posterior and the cut-posterior for $\eta$ both have closed-form expressions. We first derive these.
@@
Full Bayes Posterior.
@@
We can thus compare the marginal distribution of $\eta$ across four methods: the Full Bayes posterior, true cut-posterior, NeVI-Cut estimate of the cut-posterior, and estimated cut-posterior using the Gaussian variational approximation of Yu et al., (2023) (Gaussian VA-Cut).
We evaluate each method using 95% weighted interval score, Continuous Ranked Probability Score (CRPS; Gneiting and Raftery,, 2007), mean squared error (MSE), and 95% coverage as comparison metrics. CRPS measures the accuracy of the predictive distribution by comparing it to the observed value. The 95% weighted interval score assesses the accuracy of prediction based on the quantiles of the predicted samples. The results are presented in Table 1.
S6.2 Using Propensity Scores in Misspecified Outcome Model
In this section, we present the details of the propensity score analysis.
The propensity score model (upstream) is specified as
The true outcome model is given by:
with $\gamma^{*}=(0,0.6,0.5,0.4,0.3,0.2,0.1)$.
Appendix S7 HPV Data
Posterior samples of $\gamma_{i}$ are then passed to the downstream analysis, which models the relationship between HPV prevalence and cervical cancer incidence. Let $w=(w_{1},\ldots,w_{13})^{\top}$ denote the number of observed cervical cancer cases. The downstream analysis model specifies
Our target is the joint and marginal posterior distributions of $\eta=(\eta_{1},\eta_{2})^{\top}$. We compare the full Bayes and five different cut-posterior approaches: multiple imputation-based nested MCMC, Gaussian VA-cut (Yu et al.,, 2023), OpenBUGS-Cut which obtains cut-posteriors via the cut function in OpenBUGS with adaptive Metropolis (1D adapter), the Tempered-Cut algorithm of Plummer, (2015), and NeVI-Cut.
(a) Joint posterior of $(\eta_{1},\eta_{2})$
(b) Marginal posterior of $(\eta_{1},\eta_{2})$
Figure S2: Posterior distributions for the downstream model parameters in the HPV data analysis. The left panel shows joint posteriors of $(\eta_{1},\eta_{2})$ under full Bayes and different cut-Bayes methods. The right panel showcases marginal posteriors of $(\eta_{1},\eta_{2})$ using different cut-Bayes methods compared with multiple imputation-based nested MCMC.
The left panel of Figure S2 presents the joint posterior distributions of $\eta$ under the full Bayes and cut-Bayes models. The right panel shows pairwise comparisons of the marginal densities for the four cut-Bayes methods against the baseline cut method of nested MCMC. As noted in prior research (such as, Plummer,, 2015; Yu et al.,, 2023), the full Bayes and cut-posteriors differ substantially, underscoring the sensitivity of inference to model misspecification. The figures on the right panel show that all methods generally produce a reasonable approximation to the nested MCMC cut-posterior. This is not surprising as the cut-posterior is clearly unimodal and reasonably Gaussian-shaped. Overall, NeVI-Cut produces results that most closely align with nested MCMC. Gaussian-VA-cut and Tempered Cut tend to overestimate the spike in $\eta_{22}$, while OpenBUGS-Cut underestimates the spike in $\eta_{21}$.
Appendix S8 Estimating Cause-Specific Mortality Fractions in Mozambique
We present more scientifc context, implementation details, and additional results for the analysis of Section 6.1. In many low- and middle-income countries (LMICs), complete diagnostic autopsies are often not feasible (Nichols et al.,, 2018), particularly for children and neonatal deaths. As a result, verbal autopsies (VA), a WHO-standardized interview of the caregiver, are widely used for population-level mortality surveillance (Bassat et al.,, 2017; Menendez et al.,, 2017). VA responses are passed through pre-trained classifiers, called computer-coded verbal autopsy (CCVA) algorithms, which make individual-level causes of death (COD) predictions. These predicted CODs are then aggregated over the entire dataset to estimate the cause-specific mortality fractions (CSMFs), which are the prevalences, i.e., the fraction of deaths in the population attributable to each cause.
Extensive benchmarking against gold-standard data with more comprehensive COD ascertainment, like minimally invasive tissue sampling (MITS), has revealed that CCVA algorithms often misclassify individual-level COD. The bias due to this misclassification propagates to any downstream use of these predictions, like aggregating them to obtain CSMF estimates. To address this, a VA-calibration framework has been developed that leverages limited gold-standard data (labeled COD) to estimate the confusion matrices of the CCVA classifiers, and then uses them to correct for the misclassification bias in downstream analysis to estimate CSMFs (Datta et al.,, 2020; Fiksel et al.,, 2022, 2023; Gilbert et al.,, 2023; Pramanik et al., 2025b, ). VA-calibration has been used to calibrate VA-only data collected in the Countrywide Mortality Surveillance for Action (COMSA) program in Mozambique (COMSA-Mozambique; Macicame et al.,, 2023) and produce national-level CSMF estimates in neonates (0-27 days) and children (1-59 months).
However, the current VA-calibration implementation either models the confusion matrices and CSMFs jointly (full Bayes), or uses uncertainty-quantified estimates of confusion matrices from the upstream analysis to construct a parametric prior for the downstream CSMF estimation (sequential Bayes). Either method is essentially targeting a joint full Bayes posterior of the confusion matrices and the CSMFs. However, the downstream CSMF model is likely to be misspecified for the confusion matrices, as these confusion matrices are estimated based on limited gold-standard data collected in a clinical setting, and are being used to calibrate VA data collected in a community setting. Hence, it is undesirable to have feedback from the VA data to the estimates of confusion matrices.
In our implementation of the methods, for NeVI-Cut, following notations in Section 3.1, we consider the class of conditional distributions $q_{\vartheta}(p\,|\,\Phi)$ corresponding to the law
$p=T_{\vartheta}(\text{vec}(\Phi_{-C}),Z),\quad Z\sim N(0,I_{C}).$
[…]
@@
The results for neonates are presented in Section 6.1. We briefly summarize the children results here. Figure S3 displays the estimated densities of the cause-specific mortality fractions (CSMFs) for the child age group (1–59 months). Notably, the sequential Bayes posteriors diverge from the cut-posterior for several causes, including malaria, pneumonia, diarrhea, and other infections. Among the cut-Bayes approaches, NeVI-Cut aligns with the nested MCMC more closely, whereas the parametric cut, based on a Dirichlet variational family, shows limited flexibility and induces estimation bias.
@@
Figure S3: Posterior distributions of calibrated cause-specific mortality fractions (CSMFs) for children aged 1-59 months, based on VA-only data collected by the Countrywide Mortality Surveillance for Action (COMSA) program in Mozambique. NeVI-Cut effectively captures complex distributional characteristics, such as multimodality and skewness, achieving performance comparable to nested MCMC. Sev. mal. and Oth. Inf. denote severe malnutrition and other infections.